Find the product using suitable properties.
738 × 103

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given 783 × 103
= 783 × (100 + 3)
= 783 × 100 + 783 × 3 (using distributive property)
= 78300 + 2214
= 76014

Thanks